OpenStudy (anonymous):

determine whether 15 is the solution 5(x-1)=70 is 15 a solution?

OpenStudy (anonymous):

Ya it is. 5x-5=70 5x=75 x=15

OpenStudy (aroub):

5(15-1)=70 75-5=70 70=70 So, yes it is a solution.